# EvaGaussians: Event Stream Assisted Gaussian Splatting from Blurry Images

### v1

@@ Abstract
3D Gaussian Splatting (3D-GS) has demonstrated exceptional capabilities in 3D scene reconstruction and novel view synthesis. However, its training heavily depends on high-quality, sharp images and accurate camera poses. Fulfilling these requirements can be challenging in non-ideal real-world scenarios, where motion-blurred images are commonly encountered in high-speed moving cameras or low-light environments that require long exposure times. To address these challenges, we introduce Event Stream Assisted Gaussian Splatting (EvaGaussians), a novel approach that integrates event streams captured by an event camera to assist in reconstructing high-quality 3D-GS from blurry images. Capitalizing on the high temporal resolution and dynamic range offered by the event camera, we leverage the event streams to explicitly model the formation process of motion-blurred images and guide the deblurring reconstruction of 3D-GS. By jointly optimizing the 3D-GS parameters and recovering camera motion trajectories during the exposure time, our method can robustly facilitate the acquisition of high-fidelity novel views with intricate texture details. We comprehensively evaluated our method and compared it with previous state-of-the-art deblurring rendering methods. Both qualitative and quantitative comparisons demonstrate that our method surpasses existing techniques in restoring fine details from blurry images and producing high-fidelity novel views. Code will be released at project page.
@@ 1 Introduction
Reconstructing accurate scene-level 3D representations from 2D image collections has presented a persistent challenge within the field of computer vision and computer graphics. This task stands as a fundamental component in various vision applications, such as virtual reality [43; 11], robotics navigation [33; 51; 44], scene understanding [10; 16; 17], and many others, thereby prompting significant research efforts over the last decades. Amid pioneering works, Neural Radiance Fields (NeRFs) [23] achieves notable success in generating high-fidelity novel views by utilizing deep neural networks with a differentiable volume rendering technique [13; 21]. Despite their ability in novel view synthesis, NeRFs often suffer from poor training and rendering efficiency, limiting their applications in real-time scenarios.
@@
Recently, 3D Gaussian Splatting (3D-GS) [9] has revolutionized the field of 3D reconstruction. It extends the implicit neural representation of NeRFs to explicit 3D Gaussians with lightweight learnable parameters, and leverages a tile-based rasterization technique to render novel views, thereby surpassing NeRFs in both training and rendering efficiency as well as novel view synthesis quality. However, the optimization of 3D-GS heavily relies on accurate camera poses and point cloud initialization produced by COLMAP [35], which necessitates high-quality training images without blurring and with adequate lighting. Nevertheless, fulfilling such conditions can be challenging in real-world situations. For example, in UAVs and robotics, rapid camera movement is common when capturing images or recording videos, which often result in significant motion blur, especially in low-light conditions that require longer exposure times. The mismatched features between blurred images can lead to inaccurate pose calibrations and point clouds in COLMAP [35], or even cause it to fail in recovering camera poses, thereby hindering the training of 3D-GS.
@@
Recent studies have demonstrated the significant potential of event-based cameras in mitigating motion blur in images captured by traditional frame-based cameras [28; 8; 15]. Serving as an innovative bio-inspired visual sensor, event cameras asynchronously report the logarithmic intensity changes of each pixel captured, and can record higher temporal resolution and dynamic range data in contrast to conventional cameras. Motivated by this, prior works [30; 2], have attempted to leverage the event streams captured by event cameras to supervise the training of NeRFs. However, achieving real-time rendering and synthesizing high-fidelity novel views with intricate details poses substantial challenges for these implicit methods.
To tackle these challenges, we propose Event Stream Assisted Gaussian Splatting (EvaGaussians), which integrates the event stream recorded by event cameras into the optimization process of 3D-GS, addresses the issues stemming from motion-blurred images captured in challenging real-world scenarios. As shown in Figure. 1, capitalizing on the high temporal resolution and dynamic range offered by the event camera, we use the recorded event streams to explicitly model the formation process of motion-blurred images, and leverage a blur reconstruction loss to facilitate the initial deblurring reconstruction of 3D-GS. In addition to image-level supervision, we also employ an event reconstruction loss, which converts the rendered images into event streams using a differentiable event simulator [31; 7], and uses the real-captured high-frequency event streams as supervision, further aiding in fine detailed deblurring. By jointly optimizing the 3D-GS parameters and the camera motion trajectory of each blurry image, our method is capable of recovering a high-quality 3D-GS representation, thereby enabling generating high-fidelity novel views in real time. We comprehensively evaluate EvaGaussians on a novel synthetic dataset containing diverse scenes with various scales, and a newly collected real-world dataset recorded using a Color DAVIS346 event camera [14]. Extensive experiments demonstrate the effectiveness of our approach in both scenarios. To summarize, our contributions can be delineated as follows:
@@
  • We propose Event Stream Assisted Gaussian Splatting (EvaGaussians), the first framework tailored for reconstructing a high-quality 3D-GS from motion-blurred images with the assistance of event camera. Once trained, our method is capable of recovering intricate details of the input blurry images and allows high-fidelity real-time novel view synthesis.
  • We comprehensively evaluate the proposed method and compare it with several strong baselines, both qualitative and quantitative results demonstrate that our method achieves superior quality and outperforms previous state-of-the-art deblurring rendering methods.
@@
  • We contribute two novel datasets, including a novel synthetic dataset containing diverse scenes with various scales, and a real-world dataset captured using the Color DAVIS346 event camera [14]. We will publicly release our code and dataset for future research.
@@ 2 Related Works
2.1 Event Camera
Event-based cameras, a type of bio-inspired camera, have recently gained popularity in the field of computer vision [29; 32; 22] due to their high dynamic range and exceptional temporal resolution. It has been incorporated in the field of frame interpolation [47; 39; 26], object detection and tracking [6; 45], optical flow estimation [1; 27; 49] and so on. Several methods have been proposed to leverage the outstanding properties of event cameras to image deblurring task. To name a few, Event-based double integral (EDI) model [28] achieves model-based image deblurring by explicitly modeling the relationship between events triggered during the exposure time and the captured blurry frames. Following EDI, [8] designs a learning-based hybrid neural network that integrates both visual and temporal event information to more robustly dealing with real-world motion blur. [36] further utilizes cross-model attention to improve the deblurring quality. We recommend referring to [5] for a comprehensive understanding of event camera and its applications.
@@ 2.2 Reconstructing 3D Scene from Blurry Images
Reconstructing a high quality 3D Scene typically requires high-fidelity, sharp images as supervision. However, motion-blurred images are often occurred in real world scenarios, thus hindering accurate reconstruction of 3D scenes. Several works have been proposed to address this problem. For example, Deblur-NeRF [19] and DP-NeRF [12] attempt to learn a blur formation kernel to model the image blurring process. BAD-NeRF [41] further physically models the blurry images formation process, and adopts a bundle-adjustment strategy to jointly optimize NeRF parameters and the camera poses during the exposure time. Recently, ${\textnormal{E}^{2}}$eRF [30] and EvDeblurNeRF[3] propose to utilize event streams captured by event camera to supervise NeRF training, achieving better deblurring rendering results. However, these NeRF-based methods lack real-time rendering capabilities and suffer from extended training times. With the rapid advancement of 3D-GS, a concurrent work, BAD-Gaussians [48], proposes to utilize 3D-GS as representation and follow the blur modeling and bundle-adjustment strategy adopted in BAD-NeRF to achieve deblurring reconstruction. Although it achieves real-time rendering and faster convergence compared with prior works, it still struggles to handle severely blurred images in which COLMAP [35] will fail to produce the initial point clouds. Furthermore, it employs linear interpolation between the start and end camera poses to model the camera trajectory during exposure time, necessitating careful selection of these poses for more stable optimization.
@@ 3 Method
3.1 Preliminaries
3D Gaussian Splatting. 3D-GS [9] represents the scene with a series of sparse 3D Gaussian distributions. Each Gaussian is parameterized by an anisotropic covariance $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$ and a mean value $\mu\in\mathbb{R}^{3}$:
$\mathbf{G}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\mu)^{\top}\mathbf{\Sigma}^{-1}(\mathbf{x}-\mu)},$ (1)
where the covariance matrix $\mathbf{\Sigma}$ can be further factorized into a scaling matrix $\mathbf{S}\in\mathbb{R}^{3}$ and a rotation matrix $\mathbf{R}\in\mathrm{SO}(3)$, represented as $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top}.$ This factorization guarantees the covariance matrix’s positive semi-definiteness and simultaneously mitigates the learning complexity of 3D Gaussians. To render an image given a specific camera pose, the covariance matrix in camera coordinates, denoted as $\mathbf{\Sigma}^{\prime}$, can be calculated by applying a viewing transformation $\mathbf{W}$ [52], computed as $\mathbf{\Sigma}^{\prime}=\mathbf{J}\mathbf{W}\mathbf{\Sigma}\mathbf{W}^{\top}\mathbf{J}^{\top},$ where $\mathbf{J}$ represents the Jacobian of the affine approximation of the projective transformation, while $\mathbf{W}$ denotes the world to camera transformation matrix. Subsequently, the color of each pixel on the image plane is determined by blending Gaussians arranged in accordance with their respective depths, calculated as $\mathbf{C}=\sum_{i=1}^{N}\mathbf{T}_{i}\alpha_{i}\mathbf{c}_{i},$ where $\alpha_{i}$ symbolizes the density of the Gaussian point, which is computed by multiplying a Gaussian with covariance $\mathbf{\Sigma}$ by its corresponding opacity. Despite the exceptional capabilities of 3D-GS in 3D scene reconstruction and novel view synthesis, its training depends on high-quality sharp images and accurate camera poses, which poses substantial challenges in non-ideal real-world situations where motion-blurred images are prone to be captured.
@@ 3.1 Preliminaries
We can then represent the logarithmic intensity change as: $\mathbf{L}(t)-\mathbf{L}(f)=c\,\mathbf{E}(t)$, rewrite as $\mathbf{L}(t)=\mathbf{L}(f)+c\,\mathbf{E}(t)$, and subsequently obtain the actual intensity change:
$\mathbf{I}(t)=\mathbf{I}(f)\cdot\exp(c\,\mathbf{E}(t)).$ (3)
Therefore, when a captured image $\mathbf{I}(f)$ is given at time $f$ and the event stream is recorded during the time interval $[f,t]$, the image $\mathbf{I}(t)$ can be obtained through Eq. 3.
Figure 1: Overview of EvaGaussians. Our method seamlessly integrates the event streams captured by an event camera into the training of 3D-GS to robustly handle motion-blurred images. We adopt Event-based Double Integral (EDI) [28] for blur modeling and preprocessing, yielding initial camera trajectories and a sparse point cloud for 3D-GS training. By jointly optimizing the 3D-GS parameters and the camera motion trajectories using a blur reconstruction loss and an event reconstruction loss, our method facilitates high-quality 3D-GS reconstruction and novel view synthesis.
3.2 Modeling Motion-blurred Images Using Event Streams
The image formation process of digital cameras entails the accumulation of photons during the exposure time, which is then converted into electric charges. The motion-blurred images are resulted from camera movements during the exposure time. As introduced in [28], this phenomenon can be mathematically represented as follows:
@@ 3.2 Modeling Motion-blurred Images Using Event Streams
where $\mathbf{B}$ is the captured blurry image, which is equivalent to averaging the instantaneous latent images $\mathbf{I}(t)$ during the exposure time $[f-\tau/2,f+\tau/2]$. The optimization of 3D-GS requires camera calibration and point cloud initialization using COLMAP [35], which can fail when handling images with severe motion blur. To preliminarily obtain the camera poses and point clouds, we preprocess the motion-blurred images using the Event-based Double Integral (EDI) [28] model. Specifically, the EDI model is derived by substituting Eq. 3 into Eq. 4:
@@
Given the predefined threshold $c$, the captured blurry image $\mathbf{B}$, and the recorded event stream $\mathbf{E}(t)$, the EDI model (Eq. 5) allows the derivation of $\mathbf{I}(f)$, following which the latent image $\mathbf{I}(t)$ at any instance within the exposure time can be derived through Eq. 3. Then, as shown in Figure. 1, we uniformly sample $n$ time stamps during the exposure time and obtain a set of latent images that contains rich edge features, denoted as $\{\mathbf{I}_{i}\}_{i=0}^{n-1}$, and estimate their poses as well as obtain the initial point cloud of the scene using COLMAP [35]. Following this, a straightforward approach to optimize the 3D-GS is utilizing the preprocessed latent images as supervision. However, we found that this approach leads to poor reconstruction results, as the latent images, although providing more features than the original blurry image, still exhibits relatively low visual quality and also introduces inaccuracies into the estimated poses.
To more robustly recover a sharp 3D-GS from the motion-blurred images, we model the blurry image formation process using the EDI-produced camera poses and jointly optimize them with the 3D-GS parameters in a bundle adjustment manner [41]. Specifically, denote the estimated poses of each blurry view as $\{\mathbf{P}_{i}\}_{i=0}^{n-1}$, which roughly depicts the camera motion trajectory during the exposure time, we add each of them a learnable offset $\{\mathbf{d}_{i}\}_{i=0}^{n-1}$ as correction parameters. The resulting camera poses are defined as $\{\mathbf{\widetilde{P}}_{i}\}_{i=0}^{n-1}$, where $\mathbf{\widetilde{P}}_{i}=\mathbf{P}_{i}+\mathbf{d}_{i}$. During the optimization process, for each blurry view, we can simultaneously render $n$ images $\{\mathbf{\widetilde{I}}_{i}\}_{i=0}^{n-1}$ from the 3D-GS along the approximated camera trajectory, and simulate the formation of motion-blurred images using a discrete approximation of Eq 4, expressed as:
$\mathbf{\widetilde{B}}=\frac{1}{n}\sum_{i=0}^{n-1}\mathbf{\widetilde{I}}_{i}.$ (6)
For the total $K$ real captured blurry images $\{\mathbf{B}^{j}\}_{j=0}^{K-1}$, we can thus obtain their simulated versions $\{\mathbf{\widetilde{B}}^{j}\}_{j=0}^{K-1}$ through each learnable camera motion trajectory.
3.3 Loss Functions
Blur Reconstruction Loss. With the simulated blurry images, we use the real captured blurry images $\{\mathbf{B}^{j}\}_{j=0}^{K-1}$ to serve as image level supervision. Specifically, for each blurry image $\mathbf{B}^{j}$ and its simulated version $\mathbf{\widetilde{B}}^{j}$, we employ a blur reconstruction loss to minimize their photometric error, expressed as

### v2

@@ Abstract
3D Gaussian Splatting (3D-GS) has demonstrated exceptional capabilities in synthesizing novel views of 3D scenes. However, its training is heavily reliant on high-quality images and precise camera poses. Meeting these criteria can be challenging in non-ideal real-world conditions, where motion-blurred images frequently occur due to high-speed camera movements or low-light environments. To address these challenges, we introduce Event Stream Assisted Gaussian Splatting (EvaGaussians), a novel approach that harnesses event streams captured by event cameras to facilitate the learning of high-quality 3D-GS from blurred images. Capitalizing on the high temporal resolution and dynamic range offered by event streams, we seamlessly integrate them into the initialization and optimization of 3D-GS, thereby enhancing the acquisition of high-fidelity novel views with intricate texture details. To remedy the absence of evaluation benchmarks incorporating both event streams and RGB frames, we present two novel datasets comprising RGB frames, event streams, and corresponding camera parameters, featuring a wide variety of scenes and various camera motions. We then conduct a thorough evaluation of our method, comparing it with leading techniques on the provided benchmark. The comparison results reveal that our approach not only excels in generating high-fidelity novel views, but also offers faster training and inference speeds. Video results are available at the project page.
@@ 1 Introduction
Figure 1: EvaGaussians integrates blurry images and event streams to reconstruct sharp 3D-GS for novel view synthesis.
@@
Novel view synthesis from 2D image collections has presented a persistent challenge within the field of computer vision and computer graphics. This task stands as a fundamental component in various vision applications, such as virtual reality [45, 16, 48, 36], robotics navigation [32, 54, 46], scene understanding [14, 19, 20], and many others, thereby prompting significant research efforts over the last decades. Amid pioneering works,
@@
3D Gaussian Splatting (3D-GS) [13] achieves notable success in generating high-fidelity novel views. It learns 3D Gaussians with the lightweight learnable parameters, and leverages a tile-based rasterization technique to render novel views, thereby surpassing NeRFs [26] in both training and rendering efficiency. However, the optimization of 3D-GS heavily relies on accurate camera poses and point cloud initialization produced by COLMAP [34], which necessitates high-quality images without blurring and with adequate lighting. Fulfilling such conditions can be challenging in real-world situations. For example, in UAVs and robotics, rapid camera movement is common when capturing images or recording videos, which often result in significant motion blur. The mismatched features between blurred images can lead to inaccurate pose calibrations and point cloud initialization, thereby hindering the training process of 3D-GS.
@@
Recent studies have demonstrated the significant potential of event-based cameras in alleviating motion blur in images captured by conventional frame-based cameras [29, 11, 18, 37, 49, 12, 36]. Serving as an innovative bio-inspired visual sensor, event cameras asynchronously report the logarithmic intensity changes of each pixel captured, and can record higher temporal resolution and dynamic range data in contrast to conventional cameras. Motivated by this, prior works [30, 2], have attempted to leverage the event streams captured by event cameras to supervise the training of NeRFs. However, achieving real-time rendering and synthesizing high-fidelity novel views with intricate details poses substantial challenges for these methods.
To address these challenges, we introduce Event Stream Assisted Gaussian Splatting (EvaGaussians), which leverages the event streams captured by event cameras to enhance the learning of high-quality 3D-GS from motion-blurred images. Harnessing the exceptional temporal resolution and dynamic range offered by event streams, we use them to assist in the initialization of 3D-GS, and incorporate them to jointly optimize 3D-GS and camera trajectories of blurry images through a blur reconstruction loss and an event reconstruction loss. Due to the geometric ambiguity caused by blurry images, we further propose two event-assisted depth regularization terms to stabilize the geometry of 3D-GS. Through optimizing the 3D-GS in a progressive manner, our method can recover a high-quality 3D-GS that facilitates the real-time generation of high-fidelity novel views. To summarize, our contributions can be delineated as follows:
@@
  • We propose Event Stream Assisted Gaussian Splatting (EvaGaussians), a framework tailored for reconstructing a high-quality 3D-GS from motion-blurred images with the assistance of event camera. Once trained, our method is capable of recovering intricate details of the input blurry images and allows high-fidelity real-time novel view synthesis.
@@
  • We contribute two novel datasets, including a synthetic dataset containing diverse scenes with various scales, and a real-world dataset captured by the color DAVIS346 event camera [1], both feature various camera motions. We believe they will set a benchmark for future researches.
  • We conduct a comprehensive evaluation of the proposed method and compare it with several strong baselines. The results reveal that our approach not only excels in generating high-fidelity novel views but also provides faster training and inference speeds.
@@ 2.1 Reconstructing 3D Scene from Blurry Images
Reconstructing a high-quality 3D Scene typically requires high-fidelity, sharp images as supervision. However, motion-blurred images often occur in real world scenarios, thus hindering accurate reconstruction of 3D scenes. Several studies have been proposed to address this issue. For example, Deblur-NeRF [24] and DP-NeRF [17] attempted to learn a blur formation kernel to model the image blurring process. BAD-NeRF [41] further physically modeled the blurry images formation process, and adopted a bundle-adjustment strategy to jointly optimize NeRF parameters and the camera poses during the exposure time. These NeRF-based methods lacked real-time rendering capabilities and suffered from extended training times. With the rapid advancement of 3D-GS, a concurrent work, BAD-Gaussians [52], proposed to utilize 3D-GS as representation and follow the blur modeling and bundle-adjustment strategy adopted in [41] to achieve deblurring reconstruction. Although it achieved real-time rendering and faster convergence compared with prior works, it still struggled to handle severely blurred images in which COLMAP [34] will fail to produce the initial point clouds. Furthermore, it employed linear interpolation between the start and end camera poses to model camera trajectory during exposure time, necessitating careful selection of poses for more stable optimization.
2.2 Reconstructing 3D Scene from Event Streams
Motivated by the exceptional properties offered by event cameras, several studies attempted to reconstruct 3D scenes from event streams captured by event cameras, particularly in low-light conditions with fast camera motion. For example, EventNeRF [33], Ev-NeRF [10] and other concurrent works [43, 44, 47, 51, 48] explored the reconstruction of a 3D representation from a rapidly moving event camera. Robust e-NeRF [22] and its variants [23] further extended this task to the more challenging scenario of non-uniform camera motion, taking into account the refractory period of event cameras. These methods were typically designed to be supervised solely by information captured from a single event camera. Recently, E-NeRF [15], ${\textnormal{E}^{2}}$eRF [30], and EvDeblurNeRF [3] proposed to jointly utilize event streams captured by event cameras and motion-blurred images captured by standard frame-based cameras to reconstruct a NeRF representation. Compared to methods that rely solely on event cameras, these methods can recover accurate color details. Additionally, in contrast to RGB-only methods, they are better at handling motion blur. However, these NeRF-based methods suffer from long training and inference times, and face instability during training, which limit their further application.
Figure 2: Overview of EvaGaussians. We use event streams to assist in the initialization of 3D-GS and incorporate them to jointly optimize both 3D-GS and the camera trajectories of blurry images during the exposure time, utilizing a blur reconstruction loss and an event reconstruction loss. Additionally, we propose two event-assisted depth regularization terms to stabilize the geometry of 3D-GS.
@@ 3 Method
3.1 Preliminary
@@
We can then represent the logarithmic intensity change as: $\mathbf{L}(t)-\mathbf{L}(s)=c\,\mathbf{E}(t)$, rewrite as $\mathbf{L}(t)=\mathbf{L}(s)+c\,\mathbf{E}(t)$, and subsequently obtain the actual intensity change:
$\mathbf{I}(t)=\mathbf{I}(s)\cdot\exp(c\,\mathbf{E}(t)).$ (2)
Therefore, when an image $\mathbf{I}(s)$ is captured at time $s$, and the event stream is recorded during the time interval $[s,t]$, the image $\mathbf{I}(t)$ can be obtained by warping $\mathbf{I}(s)$ using Eq. 2.
3.2 Event-assisted Initialization
The optimization of 3D-GS requires camera calibration and point cloud initialization using COLMAP [34]. However, this process can fail when dealing with images that have significant motion blur. Motion-blurred images are resulted from camera movements during the exposure time, which can be mathematically represented as:
@@ 3.2 Event-assisted Initialization
where $\mathbf{B}$ denotes a captured blurry image, which is equivalent to averaging the instantaneous latent images $\mathbf{I}(t)$ during the exposure time $[s-\tau/2,s+\tau/2]$.
To obtain initial camera poses and point clouds for 3D-GS optimization, we first preprocess the motion-blurred images using the Event-based Double Integral (EDI) [29] model, which can be derived through substituting Eq. 2 into Eq. 3:
@@
Given the predefined threshold $c$, a blurry image $\mathbf{B}$, and the recorded event stream $\mathbf{E}(t)$, the EDI model (Eq. 4) allows the derivation of $\mathbf{I}(s)$, following which the latent image $\mathbf{I}(t)$ at any moment within the exposure time can be estimated through Eq. 2. As shown in Figure. 2(A), given a total of $K$ blurry images $\{\mathbf{B}^{j}\}_{j=1}^{K}$, for each of them, we uniformly sample $n$ time stamps during their exposure time to obtain a series of EDI-estimated latent images rich in texture features, denoted as $\{\mathbf{I}_{i}\}_{i=1}^{n}$, then obtain their poses $\{\mathbf{P}_{i}\}_{i=1}^{n}$ and the initial point cloud of the scene using COLMAP [34].
After initialization, a straightforward approach to optimizing the 3D-GS is to use the EDI-estimated latent images and poses as supervision. However, although these images provide more texture features than the original blurry image, they still do not fully recover the ideal latent image and exhibit relatively low visual quality, which also introduces inaccuracies into the camera poses, thereby leading to unsatisfactory optimization results. To more robustly recover a sharp 3D-GS from motion-blurred images, we propose to harness the advantages of event streams and seamlessly integrate them into the optimization process of 3D-GS.
3.3 Event-assisted Bundle Adjustment
As introduced in Eq. 3, during the exposure time, a motion-blurred image can be decomposed into a series of latent images along a specific camera trajectory, which can be roughly approximated by the EDI-produced camera poses $\{\mathbf{P}_{i}\}_{i=1}^{n}$ according to Eq. 4. Motivated by this, we jointly optimize these camera poses and the 3D-GS attributes in a bundle adjustment manner [41] to simultaneously recover the blur-formation camera trajectories and a sharp 3D-GS. As shown in Figure. 2(B), we add each of the EDI-produced camera poses a learnable offset $\{\mathbf{d}_{i}\}_{i=1}^{n}$ as correction parameters, resulting a learnable camera trajectory $\{\mathbf{\widetilde{P}}_{i}\}_{i=1}^{n}$, where $\mathbf{\widetilde{P}}_{i}=\mathbf{P}_{i}+\mathbf{d}_{i}$. In each training iteration, we simultaneously render $n$ images $\{\mathbf{\widetilde{I}}_{i}\}_{i=1}^{n}$ from the 3D-GS along the camera trajectory of the blurry view, and simulate the formation of motion-blurred images using a discrete approximation of Eq 3, expressed as $\mathbf{\widetilde{B}}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{\widetilde{I}}_{i}$. Consequently, for a total of $K$ real-captured blurry images $\{\mathbf{B}^{j}\}_{j=1}^{K}$, we can obtain their simulated versions $\{\mathbf{\widetilde{B}}^{j}\}_{j=1}^{K}$ through each corresponding learnable camera trajectory.
Blur Reconstruction Loss. With the simulated blurry images, we use the real captured blurry images $\{\mathbf{B}^{j}\}_{j=1}^{K}$ to serve as image level supervision. Specifically, for each blurry image $\mathbf{B}^{j}$ and its simulated version $\mathbf{\widetilde{B}}^{j}$, we employ a blur reconstruction loss to minimize their photometric error, expressed as